OpenStudy (anonymous):

3^5 sqrt (x+2)^3+3=27

OpenStudy (solomonzelman):

where does the sqrt end?

OpenStudy (anonymous):

sqrt (x+2)^3

OpenStudy (solomonzelman):

\[3^5 \sqrt{(x+2)^3}+3=27~~~~~~~this?~~~~or,~~~~3\sqrt[5]{(x+2)^3}+3=27\]

OpenStudy (anonymous):

the second one

OpenStudy (solomonzelman):

I got disconnect, sorry.

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

@hartnn

OpenStudy (solomonzelman):

Subtract 3 from both sides\[3\sqrt[5]{(x+2)^3}=24\] Divide both sides by 3\[\sqrt[5]{(x+2)^3}=8\] \[Let~~~~(x+2)=a~~~~~~~So,\]\[\sqrt[5]{a^3}=8\] (This is not it, I'll post this so that you can process is while I am doing the rest.)

hartnn (hartnn):

uh., let her try something on her own :P

OpenStudy (solomonzelman):

Sorry

OpenStudy (anonymous):

idk how to do it @hartnn !!!!!!!!

OpenStudy (anonymous):

lol

hartnn (hartnn):

thats why we are here, tell us which step u don't get, which step are you stuck on, when u try to do it on your own

OpenStudy (anonymous):

everything lol

OpenStudy (solomonzelman):

Are you good so far?

OpenStudy (anonymous):

from what your showing me yes

OpenStudy (solomonzelman):

I got disconnected again (kt's probably the latex's fault.So \[a^{3/5}=8~~~~~~~~~->~~~~~~~~~~~~~~~~a=8^{5/3}\]\[a=\sqrt[3]{8^5}~~~~~~~~->~~~~~~~a=\sqrt[3]{8^3 \times 8^2}=\sqrt[3]{(2^3)}~^5= 2^5 = 32\]